# Sublattice-selective percolation on bipartite planar lattices

### v1

@@ I Introduction
Percolation provides a remarkably simple route to nontrivial critical phenomena that has a broad range of applications in many fields of science and engineering Flory (1941); Stauffer and Aharony (2018). In its most basic formulation, one considers an infinite periodic lattice, occupying each lattice site independently with equal probability $p$. The occupied sites from contiguous clusters of connected lattice sites, which turn out to exhibit several interesting properties. In particular, an infinite, spanning (percolating) cluster of connected occupied sites emerges for sufficiently large values of $p$ beyond a particular threshold value $p_{c}$. For example, considering a two-dimensional square lattice, the site percolation threshold $p^{\mathrm{sq}}_{c}=0.592746...$ Lee (2008); Jacobsen (2015) is known from computational studies to rather high accuracy – the exact value being however unknown to date. In the vicinity of the percolation threshold, various characteristics of the cluster distribution exhibit non-trivial scaling behavior: For $p\gtrsim p_{c}$, the strength $P$ of the percolating cluster, i.e., the probability that a given site belongs to the percolating cluster, increases with $p$ by an asymptotic power law,
@@
and the mean number of sites of a finite cluster $S$ as
@@
corresponding to the order parameter susceptibility in thermal ordering transitions. The scaling exponents describe the critical behavior at the percolation transition of the above state functions and are considered universal in the sense that within a given dimension $d$ they do not depend on details of the lattice structure (e.g., square or triangluar in $d=2$) or the kind of percolation problem considered (site, bond or also in the continuum) Stauffer and Aharony (2018). For $d=2$ the exponents are known as $\beta=5/36=0.13\bar{8}$, $\nu=4/3$, and $\gamma=43/18=2.3\bar{8}$, respectively Smirnov and Werner (2001); Stauffer and Aharony (2018). Furthermore, at the percolation threshold, the infinite cluster has a fractal dimension $d_{f}=d-\beta/\nu$, e.g., $d_{f}=91/48=1.8958\bar{3}$ for the two-dimensional case.
@@
Figure 1: Illustration of the square (a) and Lieb lattice (b). Sites from sublattice $A$ and $B$ are shown in grey (white). The square lattice is shown as a checkerboard, while the nearest neighbour bonds of the Lieb lattice are shown explicitly.
@@
In physical systems the lattice sites considered above may represent objects such as atoms in a solid state crystal, an empty site being related to a defect, e.g., an impurity atom. In many condensed matter systems, the underlying lattice structure shows further characteristics, such as being bipartite, i.e., the total set of lattice sites can be decomposed into two disjoint sublattices, denoted $A$ and $B$ in the following, with nearest neighbor sites always belonging to different sublattices (cf. Fig. 1 for two examples). If, for example in a binary system, the objects occupying the $A$ and $B$ sublattice are of different species, it becomes feasible that the susceptibility to a defect differs on the $A$ and $B$ sublattice. In the context of site percolation, such a circumstance can be described by assigning two different occupation probabilities $p_{A}$ and $p_{B}$ to sites from sublattice $A$ and $B$, respectively. Such sublattice-selective percolation was introduced by Scholl and Binder in Ref. Scholl and Binder (1980), motivated by the properties of diluted magnetic compounds, with a special focus towards experimental findings in three-dimensional spinel structures. For the latter case, they performed Monte Carlo simulations to extract the phase diagram in terms of percolating vs. non-percolating regimes of parameters $(p_{A},p_{B})$ and provide numerical evidence that the critical exponents of sublattice-selective percolation are those of the universality class for conventional site percolation (where $p_{A}=p_{B}$). Ref. Scholl and Binder (1980) also considered several other lattice structures, both in three and two dimensions, focusing on the extreme case where one sublattice is fully occupied, e.g., $p_{B}=1$. In various cases, the critical value $p_{A,c}$ for $p_{A}$ can then be expressed in terms of the percolation threshold of certain related conventional percolation problems. For example, for a square lattice with $p_{A}=1$, the threshold value $p_{A,c}$ equals the site percolation threshold for a square lattice with next-nearest neighbor connectivity. Using the concept of matching lattices, this value can furthermore be shown to be equal to $0.4072531-p^{\mathrm{sq}}_{c}=$ 0.407253 … Sykes and Essam (1964). For $p_{B}=1$, the percolation threshold for $p_{A}$ thus falls below the value of $p^{\mathrm{sq}}_{c}$ – as expected, since half of the lattice is already occupied for $p_{B}=1$. Due to the equivalence between the two sublattices of the square lattice, the problem is symmetric under the exchange of $p_{A}$ and $p_{B}$ in this case. In particular, $p_{B,c}=1-p^{\mathrm{sq}}_{c}$ along the line $p_{A}=1$, and along the symmetric line $p_{A}=p_{B}$, the critical value $p^{\mathrm{sq}}_{c}$ is recovered. A quantification of the full boundary line in the $(p_{A},p_{B})$-plane of the percolating regime was however not performed in Ref. Scholl and Binder (1980) for the basic square lattice case. Later, approximate renormalization-group (RG) calculations were used to estimate this boundary line Idogaki and Uryû (1982) and to address the question of universality (the method was furthermore employed to study sublattice-selective percolation on other lattices and sublattice geometries Idogaki and Uryu (1982); Idogaki and Uryû (1983)). However, while the RG approach finds that sublattice-selective percolation is indeed controlled by the same fixed point as the uniform case, its numerical accuracy is limited. For example, in the uniform case a best estimate of 0.610 was obtained for $p^{\mathrm{sq}}_{c}$ Idogaki and Uryû (1982). Related work, using a decoupling approximation for an effective field theory of the Ising model with sublattice-selective depletion yields results of similar accuracy Idogaki et al. (1995); Ueda et al. (2000). Since those early works on sublattice-selective percolation, several methodological advances have been achieved, such as the Newman–Ziff algorithm Newman and Ziff (2000, 2001), which allows for efficient high-accuracy numerical studies of conventional lattice-based percolation problems. It is thus feasible now to also perform a systematic and accurate exploration of sublattice-selective percolation on basic planar lattices, such as the square lattice using such advanced algorithms.
@@
Here, we perform such an analysis using an adapted version of the Newman–Ziff algorithm Newman and Ziff (2000, 2001), which we detail further below. This algorithm allows for efficient Monte Carlo studies of the phase diagram, like the original algorithm does for conventional percolation problems on periodic lattices. In addition to the square lattice, we examine the case of the planar Lieb lattice, which is also bipartite, and for which conventional site-percolation has been considered recently Oliveira et al. (2021). This lattice, which is of interest in the context of flat band physics, ferrimagnetism and topological states, is a decorated square lattice and bipartite, cf. Fig. 1. However, in contrast to the square lattice, the two sublattices $A$ and $B$ of the Lieb lattice are not equivalent, and the phase diagram of sublattice-selective percolation is non-symmetric in the $(p_{A},p_{B})$-plane. Furthermore, we use finite-size scaling in order to estimate the critical exponents for sublattice-selective percolation on the square lattice. Our results support previous conclusions Scholl and Binder (1980); Idogaki and Uryû (1982) regarding the universal properties.
@@
Figure 2: Illustration of the Bethe lattice for $z=3$. Sites from sublattice $A$ and $B$ are shown in grey (white). Note that for the infinite lattice, all sites are in fact topologically equivalent.
@@ II Bipartite Newman–Ziff Algorithm
The Newman–Ziff algorithm Newman and Ziff (2000, 2001) performs Monte Carlo sampling of the percolation problem on finite lattices of linear size $L$ and with $N\propto L^{2}$ lattice sites in two dimensions. For a given system size one measures appropriate finite-size estimators $Q_{L}$ for any of the state functions $Q$, such as the quantities $P$, $\xi$ and $S$ introduced above. In the following, we consider finite systems with periodic boundaries, and a cluster is defined to be percolating if it completely wraps around in either direction. The main idea behind the Newman–Ziff algorithm can be loosely described as first calculating a given state function $Q_{L}$ in the microcanonical ensemble and then transforming to the canonical ensemble. In the microcanonical ensemble for a bipartite lattice, the state function is a function $Q_{L}(n_{A},n_{B})$ of the number of occupied sites $n_{A}$ and $n_{B}$ for each sublattice, while in the canonical ensemble it is a function of the probabilities, $Q_{L}(p_{A},p_{B})$. This is an analogy to thermodynamics, where $n_{A}$ corresponds to the energy of the system and $p_{A}$ corresponds to the temperature.
@@
Secondly, sites on the $A$ sublattice are randomly occupied according to a random permutation of the sublattice indices. Every time a site is occupied one checks each nearest neighbour one by one, and if the neighbour is occupied and belongs to a different cluster these are merged using a union–find routine, as in the original Newman–Ziff algorithm. Every time a site is newly occupied, the state functions $Q_{L}(n_{A},p_{B})$ are added to the mean $\langle Q_{L}(n_{A},p_{B})\rangle$. In Fig. 3(b), the state in which the cluster first percolates (with periodic boundary conditions) is shown. In Fig. 3(c) , the lattice is shown after adding 75 further occupied sites, and only a few small clusters besides the percolating cluster remain.
@@
Figure 3: Illustration of the sublattice-selective percolation process for an $L=32$ square lattice and periodic boundary conditions. Occupied (empty) sites are shown in grey (white), and the percolating cluster in black. Starting from a random occupation of the $B$ sublattice sites with probability $p_{B}=0.7$, the $A$ sublattice is empty in (a), filled by $n_{A}=225$ sites in (b) and $n_{A}=300$ sites in (c).
@@
In the actual implementation, the lattice is represented by an integer array of size $N$. If a site is occupied, it is part of a tree, and each tree corresponds to one cluster. A cluster of size $s$ has one root, which has an array entry of value $-s$, while the other sites of the cluster have pointers (indices of the array) which eventually lead to the root. The idea behind the aforementioned union–find algorithm is that the union operation merges two trees and the find operation returns the root of a given occupied site. So if there are two neighbouring occupied sites on the lattice, the roots can be compared via find and if the roots are different, the trees will be merged via union. Figure 4 shows an example of this procedure. The light grey site in Fig. 4(a) has been newly added, making it a tree just consisting of a root. Next, it is merged with the right cluster in Fig. 4(b), making a larger cluster of size $s=5$. In the next step in Fig. 4(c), the roots of the light grey site and the site to the left of it are compared via the find operation, and the roots are at different lattice sites. Thus, the union operation is invoked, and the root of the left tree then points to the root of the right tree and the cluster sizes are added. Now every site in the cluster returns the same root with find. There are a couple of methods to make this procedure more efficient: Firstly, the so called path compression is the idea that the find operation is most efficient when the path to the root involves as few pointers as possible. So every time the find operation follows a path from some site to the root, all the pointers on its way are set to point directly to the root – in practice this makes the trees never deeper than a few generations. Secondly, a smaller cluster is always appended to a bigger one. The find operation for the cluster that is appended takes one step longer, hence by appending the smaller cluster, the average number of steps is again minimized.
@@
Figure 4: Example of adding a newly occupied site on a square lattice. Grey sites are occupied and the light grey site is newly occupied. The roots are marked by negative numbers and the arrows symbolize pointers.
@@
In terms of computational complexity, using approaches based on alternative methods, such as the Hoshen–Kopelman algorithm Hoshen and Kopelman (1976), one would have to build all clusters for every value of $p_{A}$ from scratch, and the time complexity would be $\mathcal{O}(N_{A}^{2})$, for each realization for a given value of $p_{B}$. In contrast, the Newman–Ziff algorithm takes time of order $\mathcal{O}(N_{A}+M_{T})$ to calculate the $N_{A}$ samples of $Q_{L}(n_{A},p_{B})$ and transform from the microcanonical to the canonical ensemble, where $M_{T}$ is the number of steps needed to calculate the sum and binomial distribution in the transformation of Eq. (6). For the system sizes used here, $M_{T}$ and thus the computational time for the transformation are comparable small. For a detailed discussion of the computational advantage of the Newman–Ziff algorithm, we refer to Ref. Newman and Ziff (2001).
@@ III Square lattice
We used the bipartite Newman–Ziff algorithm algorithm to examine sublattice-selective percolation on the square lattice, and report our numerical results in this section.
@@ III.1 Percolation Threshold
We first consider the determination of the percolation threshold line for sublattice-selective percolation on the square lattice. As noted in the previous section, for this purpose we consider a set of fixed values of $p_{B}$, and then use the bipartite Newman-Ziff algorithm to obtain $Q_{L}(p_{A},p_{B})$ for essentially any value of $p_{A}$. Denoting by $R_{L}$ the probability for the existence of a percolating cluster (and estimated by the Monte Carlo mean), this state function monotonously increases from 0 to 1 smoothly upon increasing $p_{A}$ for finite $L$. This behavior is shown for two different values of $p_{B}$ in Fig. 5, considering $p_{B}=p_{c}^{\mathrm{sq}}$, and $p_{B}=1$, respectively.
@@
In the infinite system limit, this quantity – denoted $R_{\infty}(p_{A})$ – exhibits a jump at the percolation threshold and we can obtain an estimate $p_{A,c}(L)$ of $p_{A,c}$ at the specified value of $p_{B}$ from the condition $R_{L}(p_{A,c}(L))=C$, where $C\in(0,1)$ is kept fixed upon varying $L$. The speed of convergence of $p_{A,c}(L)$ towards $p_{A,c}$ varies for different choices of $C$. For conventional site percolation Newman and Ziff Newman and Ziff (2001) suggest to take $C$ equal to the exactly known value of $R_{\infty}$ at $p_{c}^{\mathrm{sq}}$, which equals $0.6904$, and find an algebraic asymptotic convergence,
@@
conjecturing that $X=2+1/\nu$, which equals $2.75$ for $d=2$, leading thus to fast convergence. From our simulations, we observe no significant change in the threshold value of $R_{\infty}$ for the sublattice-selective case (cf., e.g., the crossing point values in Fig. 5). A formal prove of this statement would however still be valuable. In any case, we always fixed $C$ to the above value in order to fit $p_{A,c}(L)$ to the finite-size scaling in Eq. (7).
In particular, we performed finite-size calculations for system sizes between $L=16$ and $L=512$, doubling consecutive values of $L$. For each fixed value of $p_{B}$, we scaled the number of Monte Carlo steps (each corresponding to an initial configurations as in Fig. 3(a), followed by successively increasing $n_{A}$) with $L^{-3/2}$, taking $10^{7}$ steps at $L=16$. This procedure ensures a similar absolute statistical uncertainty on the estimator for $R_{L}(p_{A,c}(L))$ upon varying $L$, as it does for conventional percolation Newman and Ziff (2001): For a given probability $p_{A}$ the state function $R_{L}(p_{A})$ is either 1 or 0, so the uncertainty comes from the binomial distribution,
@@
where $M$ is the number of Monte Carlo steps. As the width of the critical region decreases as $L^{-1/\nu}$, the gradient $dR_{L}/dp_{A}$ of $R_{L}(p_{A})$ in the critical region increases as $L^{1/\nu}$. Together with Eq. (8), the uncertainty on $p_{A,c}$ thus scales as $\sigma_{p_{A,c}}\sim L^{-1/\nu}\sigma_{R_{L}}\sim L^{-1/\nu}M^{-1/2}$. To keep $\sigma_{p_{A,c}}$ constant, $M$ thus has to scale with $L^{-2/\nu}$. For $d=2$, the universal value is $\nu=4/3$, and this value will also be used for this uncertainty calculation here (this has no impact on the estimation of the critical exponents as performed in the next section – there will indeed also be no evidence that $\nu\neq 4/3$). Overall, the whole procedure was repeated one hundred times for a given $p_{B}$ in order to obtain the statistical error on all other estimated state functions. For the final transformation from $n_{A}$ to $p_{A}$, we used a narrow $p_{A}$-grid with a spacing of order $10^{-4}$.
Based on the numerical data, we obtain the values of $p_{A,c}$ for several values of $p_{B}$ reported in Tab. 1. Also included in this table are the values of the exponent $X$ that we obtain based on these fits (the values of $\chi^{2}/\mathrm{d.o.f.}$ are of $O(1)$). Overall, the estimated values of $X$ exhibit some scatter, but no significant systematic deviation from the proposed value for the conventional case Newman and Ziff (2001). The numerical results for $p_{A,c}$ that we obtain for the specific cases of $p_{B}=p_{c}^{\mathrm{sq}}$ and $p_{B}=1$ are in accord with the previously reported values, ensuring the overall accuracy of our approach.

### v2

@@ I Introduction
Percolation provides a remarkably simple route to nontrivial critical phenomena, which have a broad range of applications in many fields of science and engineering Flory (1941); Stauffer and Aharony (2018). In its most basic formulation, one considers an infinite periodic lattice, occupying each lattice site independently with equal probability $p$. The occupied sites form contiguous clusters of connected lattice sites, which turn out to exhibit several interesting properties. In particular, an infinite, spanning (percolating) cluster of connected occupied sites emerges for sufficiently large values of $p$ beyond a particular threshold value $p_{c}$. For example, considering a two-dimensional square lattice, the site percolation threshold $p^{\mathrm{sq}}_{c}=0.592746...$ Lee (2008); Jacobsen (2015) is known from computational studies to rather high accuracy – the exact value being, however, unknown to date. In the vicinity of the percolation threshold, various characteristics of the cluster distribution exhibit non-trivial scaling behavior: For $p\gtrsim p_{c}$, the strength $P$ of the percolating cluster, i.e., the probability that a given site belongs to the percolating cluster, increases with $p$ by an asymptotic power law,
@@
and the mean number of sites of a finite cluster $S$ increases as
@@
corresponding to the order parameter susceptibility in thermal ordering transitions. The scaling exponents describe the critical behavior at the percolation transition of the above state functions, and they are considered universal in the sense that within a given dimension $d$ they do not depend on details of the lattice structure (e.g., square or triangular in $d=2$) or the kind of percolation problem considered (site, bond, or also in the continuum) Stauffer and Aharony (2018). For $d=2$ the exponents are known as $\beta=5/36=0.13\bar{8}$, $\nu=4/3$, and $\gamma=43/18=2.3\bar{8}$, respectively Smirnov and Werner (2001); Stauffer and Aharony (2018). Furthermore, at the percolation threshold, the infinite cluster has a fractal dimension $d_{f}=d-\beta/\nu$, e.g., $d_{f}=91/48=1.8958\bar{3}$ for the two-dimensional case.
@@
Figure 1: Illustration of the square (a) and Lieb lattice (b). Sites from sublattice $A$ and $B$ are shown in grey (white). The square lattice is shown as a checkerboard, while the nearest-neighbor bonds of the Lieb lattice are shown explicitly.
@@
In physical systems the lattice sites considered above may represent objects such as atoms in a solid state crystal, an empty site being related to a defect, e.g., an impurity atom. In many condensed-matter systems, the underlying lattice structure shows further characteristics, such as being bipartite, i.e., the total set of lattice sites can be decomposed into two disjoint sublattices, denoted $A$ and $B$ in the following, with nearest-neighbor sites always belonging to different sublattices (cf. Fig. 1 for two examples). If, for example in a binary system, the objects occupying the $A$ and $B$ sublattice are of different species, it becomes feasible that the susceptibility to a defect differs on the $A$ and $B$ sublattice. In the context of site percolation, such a circumstance can be described by assigning two different occupation probabilities $p_{A}$ and $p_{B}$ to sites from sublattices $A$ and $B$, respectively. Such sublattice-selective percolation was introduced by Scholl and Binder in Ref. Scholl and Binder (1980), motivated by the properties of diluted magnetic compounds, with a special focus towards experimental findings in three-dimensional spinel structures. For the latter case, they performed Monte Carlo simulations to extract the phase diagram in terms of percolating vs. non-percolating regimes of parameters $(p_{A},p_{B})$, and they provided numerical evidence that the critical exponents of sublattice-selective percolation are those of the universality class for conventional site percolation (where $p_{A}=p_{B}$). Reference Scholl and Binder (1980) also considered several other lattice structures, both in three and two dimensions, focusing on the extreme case in which one sublattice is fully occupied, e.g., $p_{B}=1$. In various cases, the critical value $p_{A,c}$ for $p_{A}$ can then be expressed in terms of the percolation threshold of certain related conventional percolation problems. For example, for a square lattice with $p_{B}=1$, the threshold value $p_{A,c}$ equals the site percolation threshold for a square lattice with next-nearest neighbor connectivity. Using the concept of matching lattices, this value can furthermore be shown to be equal to $0.4072531-p^{\mathrm{sq}}_{c}=$ 0.407253 … Sykes and Essam (1964). For $p_{B}=1$, the percolation threshold for $p_{A}$ thus falls below the value of $p^{\mathrm{sq}}_{c}$ – as expected, since half of the lattice is already occupied for $p_{B}=1$. Due to the equivalence between the two sublattices of the square lattice, the problem is symmetric under the exchange of $p_{A}$ and $p_{B}$ in this case. In particular, $p_{B,c}=1-p^{\mathrm{sq}}_{c}$ along the line $p_{A}=1$, and along the symmetric line $p_{A}=p_{B}$, the critical value $p^{\mathrm{sq}}_{c}$ is recovered. A quantification of the full boundary line in the $(p_{A},p_{B})$-plane of the percolating regime was however not performed in Ref. Scholl and Binder (1980) for the basic square lattice case. Later, approximate renormalization-group (RG) calculations were used to estimate this boundary line Idogaki and Uryû (1982) and to address the question of universality (the method was furthermore employed to study sublattice-selective percolation on other lattices and sublattice geometries Idogaki and Uryu (1982); Idogaki and Uryû (1983)). However, while the RG approach finds that sublattice-selective percolation is indeed controlled by the same fixed point as the uniform case, its numerical accuracy is limited. For example, in the uniform case a best estimate of 0.610 was obtained for $p^{\mathrm{sq}}_{c}$ Idogaki and Uryû (1982). Related work, using a decoupling approximation for an effective field theory of the Ising model with sublattice-selective depletion yields results of a similar accuracy Idogaki et al. (1995); Ueda et al. (2000). Since those early works on sublattice-selective percolation, several methodological advances have been achieved, such as the Newman–Ziff algorithm Newman and Ziff (2000, 2001), which allows for efficient high-accuracy numerical studies of conventional lattice-based percolation problems. Recentlty it has been shown Rosales Herrera et al. (2021) that it is now feasible to also perform a systematic and accurate exploration of sublattice-selective percolation on basic planar lattices, such as the square lattice using such advanced algorithms.
@@
Here, we perform such an analysis using an adapted version of the Newman–Ziff algorithm Newman and Ziff (2000, 2001), which we detail further below. This algorithm allows for efficient Monte Carlo studies of the phase diagram, like the original algorithm does for conventional percolation problems on periodic lattices. In addition to the square lattice, we examine the case of the planar Lieb lattice, which is also bipartite, and for which conventional site-percolation has been considered recently Oliveira et al. (2021). This lattice, which is of interest in the context of flat band physics, ferrimagnetism, and topological states, is a decorated square lattice and bipartite; cf. Fig. 1. However, in contrast to the square lattice, the two sublattices $A$ and $B$ of the Lieb lattice are not equivalent, and the phase diagram of sublattice-selective percolation is non-symmetric in the $(p_{A},p_{B})$-plane. Furthermore, we use finite-size scaling in order to estimate the critical exponents for sublattice-selective percolation on the square lattice. Our results support previous conclusions Scholl and Binder (1980); Idogaki and Uryû (1982) regarding the universal properties.
@@
Figure 2: Illustration of the Bethe lattice for $z=3$. Sites from sublattices $A$ and $B$ are shown in gray (white). Note that for the infinite lattice, all sites are in fact topologically equivalent.
@@ II Bipartite Newman–Ziff Algorithm
The Newman–Ziff algorithm Newman and Ziff (2000, 2001) performs Monte Carlo sampling of the percolation problem on finite lattices of linear size $L$ and with $N\propto L^{2}$ lattice sites in two dimensions. For a given system size, one measures appropriate finite-size estimators $Q_{L}$ for any of the state functions $Q$, such as the quantities $P$, $\xi$, and $S$ introduced above. In the following, we consider finite systems with periodic boundaries, and a cluster is defined to be percolating if it completely wraps around in either direction. The main idea behind the Newman–Ziff algorithm can be loosely described as first calculating a given state function $Q_{L}$ in the microcanonical ensemble and then transforming to the canonical ensemble. In the microcanonical ensemble for a bipartite lattice, the state function is a function $Q_{L}(n_{A},n_{B})$ of the number of occupied sites $n_{A}$ and $n_{B}$ for each sublattice, while in the canonical ensemble it is a function of the probabilities, $Q_{L}(p_{A},p_{B})$. This is an analogy to thermodynamics, where $n_{A}$ corresponds to the energy of the system and $p_{A}$ corresponds to the temperature.
@@
Secondly, sites on the $A$ sublattice are randomly occupied according to a random permutation of the sublattice indices. Every time a site is occupied, one checks each nearest neighbor one by one, and if the neighbor is occupied and belongs to a different cluster these are merged using a union–find routine, as in the original Newman–Ziff algorithm. Every time a site is newly occupied, the state functions $Q_{L}(n_{A},p_{B})$ are added to the mean $\langle Q_{L}(n_{A},p_{B})\rangle$. In Fig. 3(b), the state in which the cluster first percolates (with periodic boundary conditions) is shown. In Fig. 3(c) , the lattice is shown after adding 75 occupied sites, and only a few small clusters besides the percolating cluster remain.
@@
Figure 3: Illustration of the sublattice-selective percolation process for an $L=32$ square lattice and periodic boundary conditions. Occupied (empty) sites are shown in gray (white), and the percolating cluster is in black. Starting from a random occupation of the $B$ sublattice sites with probability $p_{B}=0.7$, the $A$ sublattice is empty in (a), filled by $n_{A}=225$ sites in (b), and $n_{A}=300$ sites in (c).
@@
In the actual implementation, the lattice is represented by an integer array of size $N$. If a site is occupied, it is part of a tree, and each tree corresponds to one cluster. A cluster of size $s$ has one root, which has an array entry of value $-s$, while the other sites of the cluster have pointers (indices of the array) which eventually lead to the root. The idea behind the aforementioned union–find algorithm is that the union operation merges two trees and the find operation returns the root of a given occupied site. So if there are two neighboring occupied sites on the lattice, the roots can be compared via find, and if the roots are different, the trees will be merged via union. Figure 4 shows an example of this procedure. The light gray site in Fig. 4(a) has been newly added, making it a tree just consisting of a root. Next, it is merged with the right cluster in Fig. 4(b), making a larger cluster of size $s=5$. In the next step in Fig. 4(c), the roots of the light gray site and the site to the left of it are compared via the find operation, and the roots are at different lattice sites. Thus, the union operation is invoked, and the root of the left tree then points to the root of the right tree and the cluster sizes are added. Now every site in the cluster returns the same root with find. There are a couple of methods to make this procedure more efficient: First, the so-called path compression is the idea that the find operation is most efficient when the path to the root involves as few pointers as possible. So every time the find operation follows a path from some site to the root, all the pointers on its way are set to point directly to the root – in practice, this makes the trees never deeper than a few generations. Secondly, a smaller cluster is always appended to a bigger one. The find operation for the cluster that is appended takes one step longer, hence by appending the smaller cluster, the average number of steps is again minimized.
@@
Figure 4: Example of adding a newly occupied site on a square lattice. Gray sites are occupied, and the light gray site is newly occupied. The roots are marked by negative numbers, and the arrows symbolize pointers.
@@
In terms of computational complexity, using approaches based on alternative methods, such as the Hoshen–Kopelman algorithm Hoshen and Kopelman (1976), one would have to build all clusters for every value of $p_{A}$ from scratch, and the time complexity would be $\mathcal{O}(N_{A}^{2})$, for each realization for a given value of $p_{B}$. In contrast, the Newman–Ziff algorithm takes time of order $\mathcal{O}(N_{A}+M_{T})$ to calculate the $N_{A}$ samples of $Q_{L}(n_{A},p_{B})$ and transform from the microcanonical to the canonical ensemble, where $M_{T}$ is the number of steps needed to calculate the sum and binomial distribution in the transformation of Eq. (6). For the system sizes used here, $M_{T}$ and thus the computational time for the transformation are comparably small. For a detailed discussion of the computational advantage of the Newman–Ziff algorithm, we refer the reader to Ref. Newman and Ziff (2001).
@@ III Square lattice
We used the bipartite Newman–Ziff algorithm algorithm to examine sublattice-selective percolation on the square lattice, and we report our numerical results in this section.
@@ III.1 Percolation threshold
We first consider the determination of the percolation threshold line for sublattice-selective percolation on the square lattice. As noted in the previous section, for this purpose we consider a set of fixed values of $p_{B}$, and then we use the bipartite Newman-Ziff algorithm to obtain $Q_{L}(p_{A},p_{B})$ for essentially any value of $p_{A}$. Denoting by $R_{L}$ the probability for the existence of a percolating cluster (and estimated by the Monte Carlo mean), this state function increases monotonously from 0 to 1 smoothly upon increasing $p_{A}$ for finite $L$. This behavior is shown for two different values of $p_{B}$ in Fig. 5, considering $p_{B}=p_{c}^{\mathrm{sq}}$, and $p_{B}=1$, respectively.
@@
In the infinite system limit, this quantity – denoted $R_{\infty}(p_{A})$ – exhibits a jump at the percolation threshold and we can obtain an estimate $p_{A,c}(L)$ of $p_{A,c}$ at the specified value of $p_{B}$ from the condition $R_{L}(p_{A,c}(L))=C$, where $C\in(0,1)$ is kept fixed upon varying $L$. The speed of convergence of $p_{A,c}(L)$ towards $p_{A,c}$ varies for different choices of $C$. For conventional site percolation, Newman and Ziff Newman and Ziff (2001) suggest taking $C$ equal to the exactly known value of $R_{\infty}$ at $p_{c}^{\mathrm{sq}}$, which equals $0.6904$, and finding an algebraic asymptotic convergence,
@@
conjecturing that $X=2+1/\nu$, which equals $2.75$ for $d=2$, leading thus to fast convergence. From our simulations, we observe no significant change in the threshold value of $R_{\infty}$ for the sublattice-selective case (cf., e.g., the crossing point values in Fig. 5). A formal proof of this statement would, however, still be valuable. In any case, we always fixed $C$ to the above value in order to fit $p_{A,c}(L)$ to the finite-size scaling in Eq. (7).
In particular, we performed finite-size calculations for system sizes between $L=16$ and $L=512$, doubling consecutive values of $L$. For each fixed value of $p_{B}$, we scaled the number of Monte Carlo steps (each corresponding to an initial configurations as in Fig. 3(a), followed by successively increasing $n_{A}$) with $L^{-3/2}$, taking $10^{7}$ steps at $L=16$. This procedure ensures a similar absolute statistical uncertainty on the estimator for $R_{L}(p_{A,c}(L))$ upon varying $L$, as it does for conventional percolation Newman and Ziff (2001): For a given probability $p_{A}$, the state function $R_{L}(p_{A})$ is either 1 or 0, so the uncertainty comes from the binomial distribution,
@@
where $M$ is the number of Monte Carlo steps. As the width of the critical region decreases as $L^{-1/\nu}$, the gradient $dR_{L}/dp_{A}$ of $R_{L}(p_{A})$ in the critical region increases as $L^{1/\nu}$. Together with Eq. (8), the uncertainty on $p_{A,c}$ thus scales as $\sigma_{p_{A,c}}\sim L^{-1/\nu}\sigma_{R_{L}}\sim L^{-1/\nu}M^{-1/2}$. To keep $\sigma_{p_{A,c}}$ constant, $M$ thus has to scale with $L^{-2/\nu}$. For $d=2$, the universal value is $\nu=4/3$, and this value will also be used for this uncertainty calculation here (this has no impact on the estimation of the critical exponents as performed in the next section – there will indeed also be no evidence that $\nu\neq 4/3$). Overall, the whole procedure was repeated 100 times for a given $p_{B}$ in order to obtain the statistical error on all other estimated state functions. For the final transformation from $n_{A}$ to $p_{A}$, we used a narrow $p_{A}$-grid with a spacing of order $10^{-4}$.
Based on the numerical data, we obtain the values of $p_{A,c}$ for several values of $p_{B}$ reported in Table 1. Also included in this table are the values of the exponent $X$ that we obtain based on these fits (the values of $\chi^{2}/\mathrm{d.o.f.}$ (degree of freedom) are of $O(1)$). Overall, the estimated values of $X$ exhibit some scatter, but no significant systematic deviation from the proposed value for the conventional case Newman and Ziff (2001). The numerical results for $p_{A,c}$ that we obtain for the specific cases of $p_{B}=p_{c}^{\mathrm{sq}}$ and $p_{B}=1$ are in accord with the previously reported values, ensuring the overall accuracy of our approach.